If the power series {summation} an xn converges uniformly in ] α ,β [ then so does {summation} an (-x)n . true or false ? Justify


1
Expert's answer
2022-05-31T12:48:36-0400

ANSWER. The statement is true in the following cases :

1) the radius of convergence of the series "\sum_{n=0}^{\infty}a_{n}x^{n}" is "\infty"

2) the convergence set of the series is the segment "[-R,R]"

3) Points "\alpha, \beta" are not end points of the set "]-R,R[" (the convergence set of the series is the "]-R,R]" or "[-R,R[, or ]-R,R[" )

If at least point one of the points "\alpha, \beta" is the end point of the set indicated in 3) , then the statement is not true.

EXPLANATION

Note , that "|a_{n}x^{n}|=|a_{n}(-x)^{n}|" . Therefore, if one of the series

"\sum_{n=0}^{\infty}a_{n}x^{n}" (1)

"\sum_{n=0}^{\infty}a_{n}(-x)^{n}" (2)

in a point "c" absolutely converges , then the other converges absolutely. Since the power series in the interval of converges "]-R,R[" converge absolutely , then the series (1) and (2) have the same radius of converges. Since the power series converges uniformly on any segment "[\alpha, \beta]\subset ]-R,R[" , then both series (1), (2) converge uniformly on "]\alpha,\beta[\subset[\alpha, \beta]" . This proves 1),2),3).

Let series (1) diverge at one of the point "-R,R" (or both)

Suppose , that "\alpha=-R" and both series (1), (2) converge uniformly on "]-R,\beta[" . There are limits at the point "(-R)" for all "n\in \N"

"\lim_{x\rightarrow(-R)^{-}} {a_{n}}x^{n}=a_{n}(-R)^{n}\, \, , \, \lim_{x\rightarrow(-R)^{-}} {a_{n}}(-x) ^{n}=a_{n}( R)^{n}\, ." By the theorem on the

limit of the sum of the power series we have the equalities

  "\lim_{x\rightarrow (-R)^{-}}\sum_{n=0}^{\infty} a_{n}x^{n}=\sum_{n=0}^{\infty} a_{n}(-R )^{n} \\\lim_{x\rightarrow (-R)^{-}}\sum_{n=0}^{\infty} a_{n}(-x) ^{n}=\sum_{n=0}^{\infty} a_{n}R^{n}"

and the series on the right converge.Therefore, the series (1) converges on the "[-R,R]" in contrast to the assumption.